OpenStudy (anonymous):

solve: 4(a+2)-2a=10+3(a-3)

OpenStudy (anonymous):

Can you distribute 4 to the brackets on left hand side??

OpenStudy (anonymous):

4a+8-2a

OpenStudy (anonymous):

Yep good.. Now : 4a - 2a = ??

OpenStudy (anonymous):

2a

OpenStudy (anonymous):

Yep so what are you left with on left hand side now ??

OpenStudy (anonymous):

..ummm do i start in the ()?

OpenStudy (anonymous):

a-3 = -3a?

OpenStudy (anonymous):

No you are simply left with 2a + 8 now on left hand side.. Now solve for Right hand side: Distribute 3 to the brackets..

OpenStudy (anonymous):

Multiply a with 3 and then multiply a with -3 and tell me what did you get ??

OpenStudy (anonymous):

. 10+3a-9?

OpenStudy (anonymous):

Yep : 10 - 9 = ??

OpenStudy (anonymous):

...1+3a

OpenStudy (anonymous):

Yep: So finally you have got : 2a + 8 = 1 + 3a Now subtract 2a from both the sides..

OpenStudy (anonymous):

..8=1+1a do i sub. 8 by 1 ...1 by 8? which ever?

OpenStudy (anonymous):

a=7

OpenStudy (anonymous):

Yes good.. You are right... Well Done..

OpenStudy (anonymous):

thanks you

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

Can you do one thing here that will help you to verify your answer if it is correct or not ??

OpenStudy (anonymous):

4(a+2)-2a=10+3(a-3) 4a+4*2-2a=10+3(a-3) 4a+8-2a=10+3(a-3) 4a-2a+8=10+3(a-3) 2a+8=10+3(a-3) 2a+8=10+3a-3*3 2a+8=10+3a-9 2a+8=3a+10-9 2a+8=3a+1 2a=3a+1-8 2a=3a-7 2a-3a=-7 -a=-7 a=7 check the answer: 4(a+2)-2a=10+3(a-3) 4(7+2)-2*7=10+3(7-3) 4(9)-14=10+3(4) 36-14=10+12 22=22 (CORRECT)